a) What is the buffer range for a buffer made with formic acid (HCOOH)? (Ka = 1.8 x 10-4)

b) What is the buffer range for a buffer made with CH3NH2 ? (Kb = 3.7 x 10-4) Be sure to think of the conjugate acid.
Answer #1

a)

use:

pKa = -log Ka

= -log (1.8*10^-4)

= 3.74

pH range of buffer is pKa-1 to pKa+1

SO, here buffer range is 2.74 to 4.74

Answer: 2.74 to 4.74

b)

use:

Ka = Kw/Kb

Kw is dissociation constant of water whose value is 1.0*10^-14 at 25 oC

Ka = (1.0*10^-14)/Kb

Ka = (1.0*10^-14)/3.7*10^-4

Ka = 2.703*10^-11

use:

pKa = -log Ka

= -log (2.703*10^-11)

= 10.57

pH range of buffer is pKa-1 to pKa+1

SO, here buffer range is 9.57 to 11.57

Answer: 9.57 to 11.57